# How to Leverage Predictive Uncertainty Estimates for Reducing Catastrophic Forgetting in Online Continual Learning

### v2

@@ 5.2. Empirical Results
The comparison among the different uncertainty strategies suggests that BI outperforms the other metrics in most cases. While the difference among the accuracy values is less pronounced, BI stands out in its effectiveness at mitigating CF. Finally, in Table 2 we evaluate our findings in the more realistic long-tailed data imbalance scenario. In both cases (artificial and real data imbalance), the results support our findings from the standard balanced setting: BI substantially reduces CF while delivering competitive or improved predictive performance.
Figure 3. Class-specific samples stored in the memory with respect to different sorting strategies.
Table 2. Comparison of last accuracy (A) and last forgetting (F) on long-tailed imbalanced datasets.
@@
6. Conclusion
In this paper, we propose to leverage estimates of predictive uncertainty for memory management. The investigation focuses on understanding the characteristics samples should have to alleviate CF. The results demonstrate that the most representative samples – i.e., the easiest-to-remember – are the most effective in mitigating CF while competitively performing in the classification task. Additionally, the newly proposed uncertainty score for memory management based on the Bregman Information shows its superiority in reducing CF compared to well-established uncertainty scores. The results on different realistic setups further support our findings and demonstrate that the use of predictive uncertainty estimates for reducing CF in online-CL is well-motivated.
Acknowledgements.

### v3

@@ 5.2 Empirical Results
Finally, in Table 3 we evaluate our findings in the more realistic long-tailed data imbalance scenario. In both cases (artificial and real data imbalance), the results support our findings from the standard balanced setting: BI substantially reduces CF while delivering competitive or improved predictive performance.
Table 3: Comparison of last accuracy (A) and last forgetting (F) on long-tailed imbalanced datasets.
@@
6 Conclusion
Given the conflicting nature of the strategies available in the literature for populating the memory, our investigation seeks – under an uncertainty lens – to a) clarify which characteristics of the samples alleviate CF in a consistent manner, and b) assist practitioners in selecting the appropriate memory management strategy. Starting from the examination of the properties and behaviours of popular uncertainty estimates, we identify that they mostly capture the irreducible aleatoric uncertainty and hypothesize that a better strategy should focus on the epistemic uncertainty instead. For this, we propose to use an uncertainty estimate based on the Bregman Information which, via a general bias-variance decomposition for strictly proper scores, identifies the variance term as an estimate for the epistemic uncertainty. Our findings indicate that the most representative samples – i.e., the easiest-to-remember or the least uncertain ones – are the most effective at mitigating CF (Figure 3) while maintaining competitive performance in classification tasks (Tables 1 and 2). Furthermore, the newly proposed BI-based uncertainty score for memory management shows its superiority in reducing CF compared to well-established uncertainty scores and other memory-based strategies (Figure 1). Due to its emphasis on epistemic uncertainty, the memory contains samples that are representative of their respective class distributions (Figure 4) while being close enough to the decision boundary, helping the learning model to distinctly shape the boundaries between the classes during training. The results on various realistic setups (Table 3) – not addressed in related work – further support our findings and demonstrate that the use of BI-based predictive uncertainty estimates for reducing CF in online-CL is appealing and well-motivated even in more challenging scenarios.
Acknowledgments
@@
Appendix A Appendix
A.1 List of Augmentations
The list of augmentations used for estimating the variance of the classification loss via TTA is reported in Figure 5.
  {minted}python transform_cands = [ CutoutAfterToTensor(args, 1, 10), CutoutAfterToTensor(args, 1, 20), v2.RandomHorizontalFlip(), v2.RandomVerticalFlip(), v2.RandomRotation(degrees=10), v2.RandomRotation(45), v2.RandomRotation(90), v2.ColorJitter(brightness=0.1), v2.RandomPerspective(), v2.RandomAffine(degrees=20, translate=(0.1, 0.3), scale=(0.5, 0.75)), v2.RandomResizedCrop(args.input_size[1:], scale=(0.8, 1.0), ratio=(0.9, 1.1), antialias=True), v2.RandomInvert() ]
Figure 5: List of the augmentations applied to the images for computing the uncertainty via TTA.
A.2 Time Complexity Analysis
Table 4 reports the average runtime per batch and the total runtime in seconds on CIFAR10. The increase in runtime from CBR (class-balanced, random) to BI (class-balanced, uncertainty-based) stems from the requirement to generate TTA images but remains fast in absolute terms. In fact, considering that in online-CL we are interested in updating the model every time a mini-batch of new data arrives, the average time for processing it remains reasonable with less than one second per batch. Since the batch size is the same for all datasets, this estimate reflects the average runtime per batch for all the considered datasets.
Table 4: Average runtime per batch and total runtime (in seconds) on CIFAR10.
Runtime per batch Total runtime
ER $\sim$ 0.22s $\sim$ 42s
CBR $\sim$ 0.31s $\sim$ 60s
BI (ours) $\sim$ 0.69s $\sim$ 132s
A.3 Memory Samples Analysis
In Figure 6, we analyse the composition and characteristics of the samples stored in the memory according to different strategies (random, top-k, and bottom-k). Figure 6(a) depicts the distribution of the confidence scores when samples are being stored in the memory. As one can see, although employing different strategies, the distributions look quite similar. Differently, in Figure 6(b), the distributions of the BI-based uncertainty scores for the same samples differ significantly across the strategies. This shows the ability of BI to capture the epistemic uncertainty of the samples irrespective of the corresponding confidence score.
(a) Distributions of confidence scores.
(b) Distributions of uncertainty values (BI).
Figure 6: Distributions of confidence scores (a) and uncertainty values (b) for samples stored in the memory according to different strategies (random, top-k, and bottom-k).
A.4 Relevance of Online-CL for AI-aided Medicine
In our experiments, we include BloodCell as an example of a realistic online-CL scenario that reflects a common challenge in AI-aided medicine: (i) classifying cells from blood smears is an important component in diagnostic hematology and (ii) its imbalanced nature reflects the common setting where different disease subtypes (that require differential treatment) occur with vastly different frequencies (class imbalance); with common diseases being common, hospitals tend to first see patients with the most common subtypes when collecting data for a ML-model and over time also patients with less common disease subtypes will present in the clinic. If the model is trained in the online-CL scenario it is not necessary to wait potentially a long time until sufficient patients of all subtypes have presented. Instead the model can be updated continually over time (online) whenever patients with a new subtype present. It is then crucial to mitigate CF e.g. via a memory based approach, to make sure all subtypes can be predicted well.
A.5 Pseudocode of online-CL with BI
Algorithm 1 Online-CL learning scheme
1:Input T: number of tasks, $bs$: batch size, $D_{t}$: task dataset.
2:Notation $\bm{\theta}$: model parameters, $\mathcal{M}$: memory buffer, $t$: current task, $b^{t}$ current batch of task $t$, $\bm{m}$: memory sample (size equal to $bs$).
3:while training not completed do
4:     $b^{t}\leftarrow$ GetNextBatch($D_{t}$) $\triangleright$ Get one batch from $t$
5:     if $t=0$ then
6:         $\bm{\theta}\leftarrow$ Train($b^{t}$)
7:     else
8:         $\bm{m}\leftarrow$ $\mathcal{M}$) $\triangleright$ Random sampling from memory
9:         $\bm{\theta}\leftarrow$ Train($b^{t}$, $\bm{m}$) $\triangleright$ Train with replay
10:     end if
11:     $\mathcal{M}\leftarrow$ $\mathcal{M}$, $b^{t}$)
12:end while
13:
14:function $\mathcal{M}$, $b^{t}$)
15:     for $c\in b^{t}$ do $\triangleright$ For each label $c$ in the current batch
16:         $\bm{m}_{c}\leftarrow$ $\mathcal{M}[y=c]$triangleright Extract samples with label $c$ from memory
17:         $\bm{a}_{c}\leftarrow$ $\bm{m}_{c}\cup b^{t}[y=c]$triangleright Get candidate samples for label $c$
18:         $\bm{u}_{c}\leftarrow$ $u_{BI}(\textsc{TTA}(\bm{a}_{c}))$triangleright Compute uncertainty with Eq. (6)
19:         $\mathcal{M}\leftarrow$ ReplaceSamples($\bm{u}_{c}$) $\triangleright$ Update $\mathcal{M}$ based on $\bm{u}_{c}$
20:     end for
21:     return updated $\mathcal{M}$
22:end function